# Beyond Prompt Content: Enhancing LLM Performance via Content-Format Integrated Prompt Optimization

### v2

@@ B.1 Query Format
Cascading_Statements
Options:
-A True, True
-B False, False
-C True, False
-D False, True
Answer: C
Highlight_Separator_Case
ANSWER > C
B.2 Prompt Renderer
@@
Appendix C Appendix: Initial Prompt
C.1 GSM8K
Prompt Renderer: Directly Joint Query Format: QA
{{Query placeholder}}
C.2 MATH500
Prompt Renderer: Directly Joint
Query Format: Question-Answer
Here are some examples:
[…]
{{Query placeholder}}
C.3 ARC-Challenge
Prompt Renderer: Directly Joint
Query Format: MultiChoice_QA
Here are some examples:
Choices:
A: It reduces the amount of carbon dioxide production
B: It reduces the production of oxygen
C: It decreases the greenhouse effect
D: It decreases pollutants in the air
Answer: B
{{Query placeholder}}
C.4 Big-Bench Classification
Prompt Renderer: Directly Joint
Query Format: Input-Output
Examples:
Input: Speaker 1: You do this often?’ Speaker 2: Its my first time.’
Output: no
{{Query placeholder}}
Appendix D Appendix: CFPO Results Analysis
D.1 In-context Examples and Text Length
Figure 5: Overview of in-context examples and text lengths for various tasks and models.
D.2 Examples of Optimal Prompt
@@
LLaMA-3.1-8B on GSM8K
**Decoding Mathematical Language in Real-World Scenarios**
**Ensuring Your Solution Fits the Scenario Perfectly**
**Examples to Illuminate the Path**
**Query**
{{query}}
LLaMA-3-8B-Instruct on MATH-500
- Examples: Here are some examples:
### Instruction:
### Response:
### Instruction:
### Response:
[…]
- Query:
{{query}}

### v3

@@ D.1.1 Query Format
D.1.2 Prompt Renderer
@@
D.2 Ablation Study of Format Generation Models
Format Generation Model Accuracy
Claude-3.5-haiku 58.91
Llama-3.3-70B-Instruct 61.71
Gemini-2.0 62.09
DeepSeek-R1 62.24
GPT-4 (CFPO) 63.38
Table 9: GSM8K Accuracy with Different LLMs for Format Generation (Using GPT-4 as content optimizer and LLaMA-3.1-8B as the target model).
To investigate the potential reliance on a specific LLM for format generation within our framework, we conducted an ablation study. This study evaluated the performance of several alternative LLMs in generating the formats used by CFPO, as shown in Table 9. We systematically replaced the original format generation LLM (GPT-4) with alternative models: Gemini-2.0, DeepSeekR1, Claude-3.5-Haiku, and Llama-3.3-70B-Instruct. In these experiments, GPT-4 was consistently used for content optimization, and LLaMA-3.1-8B remained the target model for evaluating performance on the GSM8K benchmark.
The results demonstrate that the choice of LLM for format generation has a relatively minor impact on overall accuracy. While variations in the specific generated formats are observed across different LLMs, the consistent performance suggests that CFPO’s effectiveness is not critically dependent on a single LLM. Specifically, as long as the LLM demonstrates the capability to generate variations and code based on the provided schema, the downstream performance on GSM8K remains stable. This robustness suggests that the core principles of CFPO are transferable and not tied to the idiosyncrasies of a particular LLM. This suggests CFPO is not overly sensitive to the choice of format generation model.
Appendix E Appendix: Examples of CFPO Optimal Prompt
@@
LLaMA-3.1-8B on GSM8K
**Decoding Mathematical Language in Real-World Scenarios**
**Ensuring Your Solution Fits the Scenario Perfectly**
**Examples to Illuminate the Path**
**Query**
{{query}}
LLaMA-3-8B-Instruct on MATH-500
- Examples: Here are some examples:
### Instruction:
### Response:
### Instruction:
### Response:
[…]
- Query:
{{query}}